OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 4)(x − 4) and (x + 4)(x + 4)

OpenStudy (anonymous):

\[(x + 4)(x − 4) =x^2-4^2 =x^2-16\] \[(x + 4)(x + 4) =(x + 4)^2 =x^2+2\times x \times4+4^2=x^2+8x+16\]

OpenStudy (anonymous):

Could you explain how your answer?

OpenStudy (anonymous):

you got*

OpenStudy (anonymous):

In order to find the product of (x+4)(x−4) we can use the formula \[(a+b)(a−b)=a^2-b^2\] and to find the product of \[(x+4)(x+4)=(x+4)^2\] we used the formula \[(a+b)^2= a^2+2ab+b^2\]

OpenStudy (anonymous):

Okay that makes sense

OpenStudy (anonymous):

Thank you so much